Grade 10Magical Mathematics[Interesting Approach]

A man borrows Rs 840 and agrees to repay with a total interest of Rs 240 in 12 Installments , each installment being less than the preceding one by Rs 8. what should be the first investment and the last investment ?

muthhu swami
12 Years agoGrade 10

1 Answer

Arun Kumar
12 Years ago
Hello Student,

\\840+240=1080 \\$AP whose sum is 1080 common difference 8$ \\ {12 (2a+11*8) \over 2}=1080 \\=>12(a+44)=1080 \\=>a+44=90 \\=>a=46
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty